# Loki: Low-Rank Keys for Efficient Sparse Attention

### v1

@@ 5.1 Experimental Setup
5.2 Results and Discussion
Let us begin our discussion with Figure 3, showing the perplexity (left) and downstream task (right) evaluation results for Loki on Llama2-7B/13B, Llama3-8B/70B, Mistral-7B, and Mixtral-8x22B models. We’ll focus on the Llama2-7B model, comparing pre-rotary (light green/purple) and post-rotary (dark green/purple) PCA transforms for different and values. For Llama2-7B, we see that the performance of both candidate transforms is similar and close to the full attention model. This trend is consistent across all the models except for Llama3-8B/70B and Mistral-7B, where the post-rotary PCA transform performs significantly worse than the pre-rotary one. For Llama3-8B, perplexity jumps from about 5 to over 10, a significant decline not seen with the pre-rotary transform. Mistral-7B shows a similar pattern. This is a surprising observation as one might expect the post-rotary PCA transform to perform better. We do not have a clear explanation for this observation and further investigation into RoPE embeddings is warranted to understand this behavior. Nevertheless, we see that at least one of the PCA transforms performs well for any model, and this is an important hyperparameter to tune when using our method. For subsequent results, we only show results from the better-performing transformation for each model. Comparing different settings, we see that using = 0.25 and = 0.25 (green), is better than using = 0.125 and = 0.5 (purple) for all models. These two settings balance speed and performance well, with the first being superior.
@@ 5.2 Results and Discussion
Next, we compare the performance of Loki with various baselines, using = 0.25 for all methods and = 0.25 for ours. Table 2 shows the perplexity results for Llama2-7B, Llama2-13B, Llama3-8B, and Mistral-7B. Loki’s perplexity drop is within 0.1 of full attention across all models, a threshold considered acceptable for attention mechanism approximations [35]. In contrast, H2O’s perplexity drop exceeds 0.1, nearing 0.2 for all models. Figure 4 confirms this trend in downstream task evaluation. Loki performs similarly to full attention for all models, except Llama3-8B, where the performance is notably worse, though still better than H2O. Importantly, on the challenging MMLU task, Loki degrades less than H2O.
@@
Comparing Loki with Exact-TopK, we find similar performance for Llama2-7B, Llama2-13B, and Mistral-7B. Exact-TopK represents the upper performance bound for Loki if it could perfectly select the top- tokens. To understand why Loki works well, we examined the top-k agreement between attention scores from low-dimensional keys and exact attention scores. Figure 5 shows a Jaccard similarity between the top- tokens selected by both methods across all layers and heads for Llama2-7B. For the settings: (, ) and (, ), evaluated in Figure 3, we see the Jaccard similarity is between 0.85 and 0.95, validating that the Loki is able to select the top- tokens with high accuracy.
@@
We now turn our attention to the generalizability of the PCA transformations used in our method. Figure 5 (right) shows the performance of Loki using PCA transformations derived from different calibration datasets (). For a given model, we consistently use the best-performing candidate PCA transformation (pre or post), based on the perplexity evaluation in Figure 3, for all the calibration datasets. We see that the performance of Loki is consistent across different calibration datasets, indicating that the PCA transformations used in our method are generalizable. This is an important observation as it shows that the PCA keys can be generated using a variety of calibration datasets and still achieve good performance.
@@
Analyzing Llama2-13B with Hugging Face Transformers exposed an interesting bottleneck (Figure 6, leftmost). Regardless of the attention type (vanilla or Loki), more than 80% of the inference time is consumed within the Hugging Face framework for appending key-value pairs for the latest token to the KV-Cache. This shared bottleneck minimizes the overall performance improvement of our optimizations. We hypothesize that using a more advanced inference system like vLLM [17] with efficient KV-Cache management could significantly reduce this append time, but leave that exploration for future work. To isolate the impact of our optimizations, the remaining plots in Figure 6 focus solely on the attention computation time, excluding the KV-Cache append time.
@@
In the second from left plot of Figure 6, we observe that Loki speeds up the total compute time spent in attention (sans KV-Cache appends) by nearly 40%! This is despite the fact that Loki incurs an extra matrix multiplication operation (orange, line 5 of Algorithm 1) to compute attention scores. In the other two figures, we compare the performance of Loki and vanilla attention for two prompt lengths - 2048 and 3072 and various generation lengths. For the shorter prompt length of 2048 we observe a speedup of around 35%, whereas for the longer prompt length of 3072, we observe a larger speedup of 40%. This trend is expected as larger prompts result in a bigger KV-Cache, amplifying the impact of our optimizations.
Limitations and Future Work: Our compute benchmarking shows that the cost of selecting the top- keys impacts overall speedup, and a custom kernel for this operation could enhance performance. Additionally, the memory cost of updating the KV-Cache is a bottleneck shared with HuggingFace’s implementation of vanilla attention. Using a more efficient KV-Cache management system is left for future work. Our method does not focus on reducing memory usage, and a potential future direction could involve utilizing CPU memory to store all the keys, and only transferring the top- keys to the GPU. The surprising observation that the pre-rotary PCA transforms perform better than the post-rotary ones for some models also warrants further investigation. Our key observation of the low intrinsic dimensionality of the keys opens up several avenues for future research, and further fine-tuning of Loki (like using per layer values) could lead to better results.
@@ 6 Conclusion
In conclusion, this work aims to devise an efficient sparse attention method, Loki, that does not compromise the model quality while reducing the computational complexity of attention in transformer models. We make a key observation that the key vectors in attention lie in a low-dimensional space, across different models and datasets. This insight, by itself, is interesting and can be investigated further in future work. Leveraging this insight, Loki uses attention scores computed in a lower dimensional space to rank and to select the top- tokens. It then uses the full dimensionality only for the selected tokens to compute the final attention. Our theoretical analysis shows that Loki can provide significant speedups in the attention step. To implement this efficiently, we develop optimized kernels that reduce data movement between the GPU and registers during the additional operations introduced by Loki. Our empirical evaluation shows that Loki performs better than popular approximation methods on a variety of models and tasks, with respect to preserving model quality. Finally, we show that Loki can provide speedups of up to 40% over the base attention empirically, making it a promising approach to address the computational challenges in transformer inference.
@@ Appendix A Comprehensive Dimensionality Analysis
A.1 Ranks Analysis for All Models
In this section, we present our dimensionality analysis results (from 3) for all the models we experimented with. Figure 7 shows the for all the models mentioned in Section 3. We observe that our findings around the low dimensionality of the keys are consistent across all models and datasets. The results for the other models are similar to the ones shown in Figure 2. The set of models we experimented with encompasses a wide range of model sizes, architecture classes (dense vs MoE models), older and newer models, and models trained on different datasets. Even with these variations, our key observation holds true. An interesting trend we observe is that the varies across layers for different models. This indicates that the intrinsic dimensionality of the keys is not uniform across layers of a model. A possible future direction could be to investigate the reasons behind this variation, and whether having a per layer in Loki could lead to better results.
Figure 8 shows the normalized eigenvalues of the covariance matrix of the keys for a few layers and heads of Llama2-7B, Mistral-7B, and Pythia-6.9B on the WikiText-2 dataset as an example. The results for the other models are similar to the ones shown here.
@@ Appendix B Comprehensive Evaluation Results
B.1 Performance of Loki on Perplexity and Downstream Tasks
@@
In this section, we present the detailed evaluation of our method on a wide range of models and tasks. Figure 11 shows the performance of Loki on perplexity and downstream tasks compared to the full attention baseline. We show the results for both pre-rotary (top) and post-rotary (bottom) PCA transformation. The models evaluated are Llama2-7B, Llama2-13B, Llama2-70B, Llama3-8B, Llama3-70B, TinyLlama-1.1B, Mistral-7B, and Mixtral-8x7B. We evaluate the models with different configurations of and for Loki. We can see that as and decrease, the performance of the model deteriorates. This is especially true when and are set to 0.125. We notice that the impact of on performance is more significant than . This is evident from the fact that performs significantly worse than for almost all the models. The two settings with and perform relatively well across all models. These settings provide a good trade-off between performance and model accuracy, with a theoretical speedup of 2.6x for both settings. All settings with preserve model quality much better but do not provide a significant speedup empirically. Table 3 and Table 4 show the same results in tabular form.
Appendix C Comparison of our top-k kernels with SparQ
As mentioned in Section 4.3, we create optimized kernels in Triton to efficiently compute the three matrix multiplications in Loki (lines 5, 8, and 9 of Algorithm 1) without creating temporary dense copies of subsets of the KV-cache. Initially, we planned to use the implementations developed by the authors of SparQ [25]. However, we discovered two major issues with their kernels. Let’s say you are multiplying two matrices of sizes and , then SparQ kernels parallelize compute along only the m dimension. However, it is well known that one can parallelize matrix multiplications along the n dimension as well and gain more performance. Thus, we add this extra dimension of parallelism to their triton kernel. Second, their kernels cannot handle non-powers of 2 number of tokens in the KV-cache, a setting which is commonly encountered in inference since we generated keys and values one at a time. Therefore, we extend their kernels to handle non-powers of two number of tokens in the KV-cache successfully. In Figure 12, we compare the performance of our kernel with sparq and vanilla PyTorch based attention for an attention layer in Llama2-7B for various sizes of the KV-cache ranging from 512 to 4096. We do this for the matmul operation of query and keys with top-k as 0.25.
@@ Appendix D Generation Examples
Appendix E PCAAttn: Using reduced-dimensional attention scores directly in transformer models
One other approach we tried is to directly use the formulation in 4.1 to compute the final attention scores. More specifically, we compute the PCA transformed query and key vectors, projected onto the first principal components, and then compute the attention scores. We only store the reduced dimension key vectors in the KV cache. We call this method PCAAttn (Algorithm 2).
Compute and Memory Analysis: When computing attention between a single query and the key vectors , the matrix multiplication has a complexity of . Using PCAAttn, the key and query vectors are reduced to dimensions and the complexity of the matrix multiplication is reduced to . Thus, we can get a speedup of in the attention dot product computation. The PCA transformation of the query and key vector generated at each step has a complexity of , which is small when . The KV-Cache memory requirement is reduced by a factor of because we only reduce the key vectors to dimensions and not the values. Additionally, the PCA adds a significantly small memory overhead of . Table 5 shows the explanation of key-budget and dimensionality for PCAAttn, along with the speedup and memory savings.
| Method | Budget | Dim. | Description | Speedup | Memory Savings |
| PCAAttn | Full | of full dimensionality used to store keys and compute attention output |

### v2

@@ 5 Experimental Setup
6 Results
We now present the comparisons of Loki with full attention and other sparse attention methods, including a comparison of the computation times.
6.1 Comparison with Full Attention
Let us begin our discussion with Figure 3, showing the perplexity (left) and short-context downstream task evaluation (right) results for Loki on different models. We focus on the Llama2-7B model, comparing pre-rotary (light green/purple) and post-rotary (dark green/purple) PCA transforms for different and values. For Llama2-7B, we see that the performance of both candidate transforms is similar. This trend is consistent across all the models except for Llama3-8B/70B and Mistral-7B, where the post-rotary PCA transform performs significantly worse than the pre-rotary one. For Llama3-8B, perplexity jumps from about 5 for the full attention to over 10, a significant decline not seen with the pre-rotary transform. Mistral-7B shows a similar pattern. This is a surprising observation since attention scores are calculated from post-rotary keys in the original attention mechanism. A possible explanation is that post-rotary PCA captures token distributions tied to specific positions in the calibration dataset, while pre-rotary PCA may generalize better by using less positional information. Nevertheless, at least one of the PCA transformations performs well for every model. For subsequent results, we only show the better-performing transformation for each model.
Figure 4 shows the performance of Loki on the LongBench tasks for the Llama2-7B-Chat model. We see that for all tasks, either one of the two candidate transforms performs similarly to full attention. For Summarization, Few Shot Learning, Synthetic, and Code Completion task categories, the best performing Loki configuration is at par or better than the full attention model. For the Single-Doc QA and Multi-Doc QA task categories, Loki performs slightly worse than the full attention model, with the biggest drop in performance observed for HotpotQA of around 3%. Comparing different settings, we see that using = 0.25 and = 0.25 (green), is better than using = 0.125 and = 0.5 (purple) for all models and tasks (short-context and long-context). These two settings balance speed and performance well, with the first being superior for accuracy.
6.2 Comparison with Other Sparse Attention Methods
Next, we compare the performance of Loki with other methods, using = 0.25 for all methods and = 0.25 for ours. Table 2 shows the perplexity results for Llama2-7B/13B, Llama3-8B, and Mistral-7B. Loki’s perplexity drop is within 0.1 of full attention across all models, a threshold considered acceptable for attention mechanism approximations [37]. In contrast, H2O’s perplexity drop nears 0.2 for all models. Figure 5 confirms this trend on short-context evaluation. Loki performs similar to full attention for all models, except Llama3-8B, where the performance is notably worse, but still better than H2O. Importantly, on the challenging MMLU task, Loki degrades less than H2O.
@@ 6.2 Comparison with Other Sparse Attention Methods
It is important to note here that Loki is designed to be compatible with other sparse attention methods. For instance, token-eviction methods like H2O delete tokens to save KV-cache memory, whereas Loki reduces memory bandwidth by selecting the top- tokens without deletion, making them orthogonal. A combined approach could involve using H2O to delete tokens, then applying Loki to select top- tokens from the remaining cache. Similarly, Loki is theoretically orthogonal to quantization methods.
@@
Comparing Loki with Exact-TopK, we find similar performance for Llama2-7B/13B and Mistral-7B. Exact-TopK represents the upper performance bound for Loki if it could perfectly select the top- tokens. To understand why Loki works well, we examined the top- agreement between Loki’s reduced dimensional attention scores and exact attention scores. Figure 6 shows the Jaccard similarity between the top- tokens selected by both methods across all layers and heads for Llama2-7B. For the settings: (, ) and (, ), evaluated in Figure 3, the Jaccard similarity is around 0.9, validating that the Loki is able to select the top- tokens with high accuracy.
6.3 Generalizability
@@
We now turn our attention to the generalizability of the PCA transformations used in our method. Figure 6 (middle) shows the performance of Loki using PCA transformations derived from different calibration datasets (). We see that the performance of Loki is consistent across different calibration datasets, indicating that the PCA transformations used in our method are generalizable. This is an important observation as it shows that the PCA keys can be generated using a variety of calibration datasets and still achieve good performance.
6.4 Computational Efficiency
@@
We now turn our attention to the computational efficiency of Loki. Analyzing Llama2-13B with Hugging Face Transformers exposed an interesting bottleneck (Figure 6, rightmost). Regardless of the attention type (vanilla or Loki), more than 80% of the time is consumed within the Hugging Face framework for appending key-value pairs of the latest token to the KV-cache. This shared bottleneck minimizes the overall performance improvement of our optimizations. We hypothesize that using a more advanced inference system like vLLM [18] could significantly reduce this append time, but leave that exploration for future work. To isolate the impact of our optimizations, the plots in Figure 7 focus solely on the attention computation time, excluding the KV-cache append time.
@@
In the left plot of Figure 7, we see that Loki speeds up the total attention compute time (excluding KV-cache appends) compared to vanilla attention across various prompt and generation lengths. For a prompt length of 3072 and generation length of 512, Loki achieves nearly a 45% speedup, despite the fact that it incurs an extra matrix multiplication operation. The breakdowns also show that the top- operation is nearly as expensive as the smaller matrix multiplications, which is a significant bottleneck. Replacing PyTorch’s top- with a custom kernel could improve this. For the shorter prompt length of 2048 we observe a speedup of around 40% (generation length = 512), slightly lower than the speedup at 3072. This trend is expected as larger prompts result in a bigger KV-cache, amplifying the impact of our optimizations.
Figure 7 (Right) shows the accuracy vs. attention time trade-off across various settings of Loki, with accuracy measured on LongBench and attention times from our microbenchmark. The previously evaluated settings, and , provide a good balance between performance and accuracy, with favoring accuracy slightly and favoring performance.
@@ 7 Conclusion
In conclusion, we introduced Loki, an algorithm for efficient sparse attention that does not compromise the model quality while reducing the computational complexity of self attention. We made a crucial observation that key vectors in attention lie in a low-dimensional space, across different models and datasets. Leveraging this insight, Loki uses attention scores computed in a lower-dimensional space to rank and select the top- most relevant tokens from the KV-cache. It then uses the full dimensionality only for the selected tokens to compute the final attention. Our theoretical analysis shows that Loki can provide significant speedups in the attention step. To implement this efficiently, we develop optimized kernels for the various sparse matrix multiplications in our approach. Our empirical evaluation shows that Loki performs better than popular approximation methods on a variety of models and tasks, with respect to preserving model quality. Finally, we show that Loki can provide speedups of up to 45% over the vanilla attention empirically, making it a promising approach to address the computational challenges in transformer inference.
Limitations and Future Work: Loki does not focus on reducing memory usage of the KV-cache currently. As mentioned previously in 6.2, it can potentially be combined with other sparse attention method for improved memory-performance-accuracy trade-offs. Another direction involves storing the KV-cache in CPU memory and transferring only the top- keys and values to the GPU [19].
While Loki outperforms vanilla attention in our benchmarks, practical deployment would require integration with efficient attention kernels like FlashAttention [7]. As seen in our compute benchmarking, the top- selection operation could introduce a bottleneck towards achieving this. Investigating this bottleneck and integrating Loki with optimized attention kernels is left for future work.
Our finding of the keys’ low intrinsic dimensionality suggests promising research directions. The variation of this dimensionality across heads and layers could further be explored. We briefly experimented with a variable policy per layer (see Appendix B.2), but did not observe significant significant improvements. A more sophisticated policy could be explored in future work.
Acknowledgments and Disclosure of Funding
This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award DDR-ERCAP0029894. Soheil Feizi was supported in part by the following grants: NSF CAREER AWARD 1942230, ONR YIP award N00014-22-1-2271, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, NSF award CCF2212458, NSF award 2229885, an Amazon Research Award and an award from Capital One.
@@ Appendix A Comprehensive Dimensionality Analysis
A.1 Complete Key-Dimensionaltiy Analysis for All Models
In this section, we present our extended dimensionality analysis results (from 3) for all the models we experimented with. Figure 8 displays the values for all models referenced in Section 3. Our analysis reveals that the low dimensionality of the keys is consistently observed across all models and datasets. Results for all models resemble those shown in Figure 2 of the main text. The models we tested cover a wide range of sizes, architecture types (dense vs. MoE), as well as older and newer architectures trained on various datasets. Despite these differences, our main observation remains robust.
An intriguing trend is the variation in across layers for different models, indicating that the intrinsic dimensionality of the keys is not uniform across model layers. A potential future direction could be to investigate the reasons for this variation from a semantic perspective.
For a more fine-grained analysis, we plot the normalized eigenvalues of the covariance matrix of the keys for a few layers and heads of Llama2-7B, Mistral-7B, and Pythia-6.9B on the WikiText-2 dataset as an example in Figure 9. Here again, we observe that the explained variance significantly decreases after the initial principal dimensions. The results for the other models are similar to the ones shown here.
@@ A.2 Variation of Rank across Attention Heads
In this section, we examine the variation in the rank at which 90% of the variance is explained () across different heads within a layer for two models: Llama2-7B and Mistral-7B. Figure 10 shows the heatmap of for the pre-rotary (top) and post-rotary (bottom) keys across all layers and heads for Mistral-7B. We observe that the is significantly lower for pre-rotary keys compared to post-rotary keys. Focusing on the pre-rotary keys, it is evident that the initial layers exhibit lower rank values than the later layers. In each layer, some heads demonstrate high-rank values, even though the median rank remains low. This suggests that certain heads in those layers are more important and leverage more complex information about the keys. Interestingly, for the post-rotary keys, we notice a pattern in which 4 out of the 8 heads in each layer share the same rank. This phenomenon may be related to how rotary embeddings are applied in Mistral-7B, as we do not observe this pattern in Llama2-7B. Further investigation is needed to understand this trend.
Figure 11 illustrates the heatmap of for the pre-rotary (left) and post-rotary (right) keys across all layers and heads for Llama2-7B. A similar trend emerges as seen in Mistral-7B, where the initial layers have lower rank values compared to the later layers. However, the same pattern in post-rotary keys that was observed in Mistral-7B is absent here, suggesting that rotary embeddings may be applied differently in Llama2-7B compared to Mistral-7B.
For both the models, we can see that in each layer, there are some heads with very high-rank values, even when the median rank is low. This might indicate that some heads in that layer are more important and use more complex information about the keys. Analysis into chosing the best reduced dimensionality based upon the the distribution of ranks across heads could be a potential future direction.
A.3 Dimensionality Analysis for Queries and Values
While our main focus has been on the dimensionality of the keys, we also performed exploratory analysis on the dimensionality of the queries and values. Figures 12 and 13 show the for the queries and values, respectively, for Llama2-7B and Llama3-70B. We observe that the queries and values also exhibit low dimensionality across all layers and heads, similar to the keys, while values tend to have a considerably higher dimensionality and close to the full dimensionality of the value vectors. This observation can intuitively be explained by the fact that both keys and queries are used to compute the scalar attention scores and thus do not need to be high-dimensional, while values are weighted by these scores and used to compute the final output, and thus need to be high-dimensional to capture the complexity of the data.
@@ Appendix B Comprehensive Evaluation Results
B.1 Performance of Loki on Perplexity and Short-Context Downstream Tasks
@@
In this section, we provide a detailed evaluation of our method across a wide range of models and tasks. Figure 14 illustrates the performance of Loki on perplexity and downstream tasks compared to the full attention baseline. We present results for both pre-rotary and post-rotary PCA transformations. The models evaluated include Llama2-7B, Llama2-13B, Llama2-70B, Llama3-8B, Llama3-70B, TinyLlama-1.1B, Mistral-7B, and Mixtral-8x7B. We assess the models using various configurations of and for Loki.
As and decrease, the model’s performance deteriorates, particularly when both are set to 0.125. Notably, the impact of on performance is more pronounced than that of . This is evident as and significantly underperform compared to and across nearly all models. The configurations of and , along with and , demonstrate relatively strong performance across all models, striking a favorable balance between performance and accuracy, with a theoretical speedup of 2.6x for both configurations. Settings with maintain model quality much more effectively but do not yield a significant empirical speedup.
B.2 Variable Dimensionality Analysis
In this section, we present our experiments utilizing a variable policy per layer in Loki. For this experiment, we set based on the per-layer explained variance, varying from 0.5 to 0.8, and plotted the average short-context task accuracies (refer to Section 5) against the compression ratio, calculated as follows:
Figure 15 presents these plots for the Llama2-7B, Llama2-13B, and Llama3-8B models. We observe that the variable policy does not yield significant improvements over the fixed policy used in our experiments. It is possible that different layers may require distinct explained variance thresholds for optimal performance, indicating that further tuning is necessary. Nevertheless, the simplicity of the fixed policy, combined with its comparable performance to the variable policy, makes it a practical choice for Loki.
Appendix C Comparison of our kernels with SparQ
As mentioned in Section 4.3, we create optimized kernels in Triton to efficiently compute the three matrix multiplications in Loki (lines 5, 8, and 9 of Algorithm 1) without creating temporary dense copies of subsets of the KV-cache. Initially, we planned to use the implementations developed by the authors of SparQ [26]. However, we discovered two major issues with their kernels. Let’s say you are multiplying two matrices of sizes and , then SparQ kernels parallelize compute along only the m dimension. However, it is well known that one can parallelize matrix multiplications along the n dimension as well and gain more performance. Thus, we add this extra dimension of parallelism to their triton kernel. Second, their kernels cannot handle non-powers of 2 number of tokens in the KV-cache, a setting which is commonly encountered in inference since we generated keys and values one at a time. Therefore, we extend their kernels to handle non-powers of two number of tokens in the KV-cache successfully. In Figure 16, we compare the performance of our kernel with SparQand vanilla PyTorch based attention for an attention layer in Llama2-7B for various sizes of the KV-cache ranging from 512 to 4096. We do this for the matmul operation of query and keys with top- as 0.25.
@@ Appendix D Generation Examples
Appendix E PCAAttn: Directly Leveraging Reduced-Dimensional Attention Scores
One other approach we tried is to directly use the formulation in 4.1 to compute the final attention scores. More specifically, we compute the PCA transformed query and key vectors, projected onto the first principal components, and then compute the attention scores. We only store the reduced dimension key vectors in the KV-cache. We call this method PCAAttn (Algorithm 2).
Compute and Memory Analysis: When computing attention between a single query and the key vectors , the matrix multiplication has a complexity of . Using PCAAttn, the key and query vectors are reduced to dimensions and the complexity of the matrix multiplication is reduced to . Thus, we can get a speedup of in the attention dot product computation. The PCA transformation of the query and key vector generated at each step has a complexity of , which is small when . The KV-cache memory requirement is reduced by a factor of because we only reduce the key vectors to dimensions and not the values. Additionally, the PCA adds a significantly small memory overhead of .